# Controlling Ensemble Variance in Diffusion Models: An Application for Reanalyses Downscaling

### v1

@@ 3.2 Controlling the variance
The generative process for DDIM involves a sequence of denoising operations alternated with the re-injection of the same noise. Our findings reveal that the key aspect of controlling the variance is selecting the number of diffusion steps. In this section, we analyze the evolution of the variance through the iterative diffusion process and we provide an intuition on how the number of diffusion steps $N$ influences the variance of the generated data. We remark that the diffusion process in DDIM is defined by sampling $\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ and then iterating through:
@@
where we defined $c_{t-\Delta t}$ as:
@@
to simplify the notation. We also recall that the diffusion schedule $\{\alpha_{t}\}_{t=0,\dots,T}$ is selected such that $\alpha_{0}\approx 1$ and $\alpha_{T}\approx 0$. Note that $\alpha_{T}\neq 0$ as we clamped its value to avoid instabilities during training, as we already discussed in section 3.1.
Let $\boldsymbol{x}_{0}$ be an $n$-dimensional random variable with Gaussian distribution such that $\mathbb{E}[\boldsymbol{x}_{0}]=\boldsymbol{0}$ and $Var(\boldsymbol{x}_{0})=\boldsymbol{I}$, and let $\{\boldsymbol{x}_{t}\}_{t=0,\dots,T}$ be the stochastic process defined in (8). Since the full variance matrix $Var(\boldsymbol{x}_{t})$ has dimension $n\times n$, which is prohibitive from a computational point of view, in this work we limit our analysis to the element-wise variance, defined as:
$\boldsymbol{v}(\boldsymbol{x}_{t})_{i}:=Var(\boldsymbol{x}_{t})_{i,i},$ (10)
which is a vector of dimension $n$, representing the variance between each pixel of $\boldsymbol{x}_{t}$ with itself, ignoring the relationship with the adjacent pixels. Note that this choice is coherent with our task, as we want to study the pixel-wise uncertainty of DDIM downscaling, and taking into consideration the correlation between adjacent pixels is beyond the scope of this analysis.
It is not difficult to show (see Appendix A) that the element-wise variance has similar properties to the classical variance, and in particular:
  1. (P1)
    $\boldsymbol{v}(a\boldsymbol{x}_{t})=a^{2}\boldsymbol{v}(\boldsymbol{x}_{t})$,
  2. (P2)
    $\boldsymbol{v}(\boldsymbol{x}_{t}+\boldsymbol{y}_{t})=\boldsymbol{v}(\boldsymbol{x}_{t})+\boldsymbol{v}(\boldsymbol{y}_{t})+2\boldsymbol{v}(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$,
where $\boldsymbol{v}(\boldsymbol{x}_{t},\boldsymbol{y}_{t})$ is the element-wise covariance between $\boldsymbol{x}_{t}$ and $\boldsymbol{y}_{t}$.
@@
$\boldsymbol{F}_{t}:=\frac{\alpha_{t}}{\alpha_{t-\Delta t}}\boldsymbol{I}+\boldsymbol{J}_{t},$
$\boldsymbol{J}_{t}:=2\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}c_{t-\Delta t}diag\left(\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})\right),$
$\boldsymbol{g}_{t}:=c_{t-\Delta t}^{2}\boldsymbol{1}.$
@@
See Appendix A. ∎
Note that (11) represents a first-order linear recurrence equation of dimension $n$ with non-constant coefficients, for which there exists a closed formula that can be used to compute $\boldsymbol{v}_{T}$, the variance of the generated data, as summarized in the following Theorem.
@@
Let $\boldsymbol{F}_{t}$, $\boldsymbol{J}_{t}$, and $\boldsymbol{g}_{t}$ defined as in Proposition 3.1, and let $\Delta t$ be a fixed step-size for DDIM, so that any timestep $t$ can be written as $i\Delta t$ for $i=0,\dots,N$. Note that $\boldsymbol{v}_{0}=\boldsymbol{1}$, the vector of all ones, as $\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$. Then, defining $\boldsymbol{F}_{(N+1)\Delta t}$ to be the identity matrix to simplify the notation, it holds:
$\boldsymbol{v}_{T}\approx\left(\prod_{i=1}^{N}\boldsymbol{F}_{i\Delta t}\right)\boldsymbol{1}+\sum_{i=1}^{N}\left[\left(\prod_{k=i+1}^{N+1}\boldsymbol{F}_{k\Delta t}\right)\boldsymbol{g}_{i\Delta t}\right].$ (12)
@@
See Appendix A. ∎
Notably, Theorem 3.2 not only allows for the explicit computation of the element-wise variance of $\boldsymbol{x}_{T}$, but it also underlines the dependence of $\boldsymbol{v}_{T}$ on the number of diffusion timesteps. Indeed, if, for example, both $\boldsymbol{F}_{i\Delta t}\geq\boldsymbol{0}$ and $\boldsymbol{g}_{i\Delta t}\geq\boldsymbol{0}$, then Equation 12 clearly implies that $\boldsymbol{v}_{T}$ monotonically grows as a function of $N$. Specifically, increasing the step size $\Delta t$ (and consequently reducing the number of timesteps $N$) results in fewer terms being included in the summation that computes $\boldsymbol{v}_{T}$, thereby decreasing its value.
@@ 4.2 Training on the full domain
The general structure of our diffusion model follows the framework established in [7]. The model utilizes a residual U-Net [1, 49] as a denoiser. Training is conducted by uniformly sampling signal-to-noise ratios from the sinusoidal noise scheduling functions, optimizing the mean absolute error (MAE) between the sampled noise and the predicted noise.
@@
Recent findings [47, 48] have demonstrated that cosine noise schedules can reduce the effective noise added by the model as image size increases. This occurs because redundancy in data, such as correlations among nearby pixels, typically increases with resolution, and the independently added noise becomes easier to denoise in larger images. Consequently, noise schedules optimized for smaller resolutions may not perform as effectively at higher resolutions. To address this issue, we introduce a constant scaling factor $\lambda>1$ to reduce the signal rate in the noise schedule. For an input $\boldsymbol{x}_{T}\sim p_{gt}(\boldsymbol{x}_{T})$ and a scaling factor $\lambda$, the noise schedule eq. 1 is modified as follows:
@@
While the core U-Net architecture remained largely unchanged, scaling to larger domains required increasing the number of residual blocks in the bottleneck. This adjustment improved performance when generating larger images [48]. The model’s performance across the increased domain remained comparable to that observed at smaller resolutions, with the bottleneck modification enhancing its ability to process higher-resolution inputs effectively.
The diffusion model is trained on the ERA5 to CERRA task for wind speed on the full domain, utilizing data from 1985 to 2010 for training. The training is run on three A100 GPUs, utilizing the AdamW optimizer.
@@ 5 Experiments
Our primary experimental objective is to evaluate the behavior of ensemble variance while changing the timestep $\Delta t$. While a comprehensive analysis of the ensemble diffusion model’s performance across the entire CERRA domain is beyond the scope of this paper, we ensure that the model operates within a realm of satisfactory performance. To achieve this, we begin—similar to [7]—by comparing our model’s performance against bilinearly interpolated ERA5 data and a residual U-Net model trained for the same task. The results, obtained by computing the Mean Squared Error (MSE) and SSIM between the high-resolution output produced by the model and the original CERRA, are shown in Table 1.
@@
[…]
SSIM ()($\uparrow$)( ↑ ) $0.751$ $0.892$ $0.923$

### v2

@@ 3.2 Controlling the variance
The generative process in DDIM consists of a sequence of denoising operations alternated with the re-injection of the same noise. Our findings reveal that the key factor in controlling the variance lies in the choice of the number of diffusion steps. In this section, we analyze the evolution of the variance throughout the iterative diffusion process and provide an intuition of how the number of diffusion steps $N$ influences the variance of the generated data. Recall that the diffusion process in DDIM is defined by sampling $\boldsymbol{x}_{0}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I})$ and then iterating through:
@@
where we define $c_{t-\Delta t}$ as:
@@
to simplify the notation. The diffusion schedule $\{\alpha_{t}\}_{t=0,\dots,T}$ is chosen such that $\alpha_{0}\approx 1$ and $\alpha_{T}\approx 0$. Note that $\alpha_{T}\neq 0$, as its value is clamped to avoid instabilities during training, as already discussed in section 3.1.
Let $\boldsymbol{x}_{0}$ be an $n$-dimensional random variable with Gaussian distribution such that $\mathbb{E}[\boldsymbol{x}_{0}]=\boldsymbol{0}$ and $Var(\boldsymbol{x}_{0})=\boldsymbol{I}$, and let $\{\boldsymbol{x}_{t}\}_{t=0,\dots,T}$ be the stochastic process defined in (8). Since the full covariance matrix $Var(\boldsymbol{x}_{t})$ has dimension $n\times n$, its computation is prohibitive for high-dimensional data such as images. Therefore, in this work we restrict our analysis to the total variance, defined as the sum of the diagonal elements of the covariance matrix:
$\boldsymbol{v}(\boldsymbol{x}_{t}):=\operatorname{Tr}\!\big(Var(\boldsymbol{x}_{t})\big)=\sum_{i=1}^{n}Var(\boldsymbol{x}_{t})_{i,i}.$ (10)
The total variance provides a scalar measure of the overall uncertainty of $\boldsymbol{x}_{t}$, summarizing how much the process spreads around its mean across all pixels while ignoring pairwise correlations. This simplification allows us to capture the global evolution of uncertainty along the diffusion trajectory, which is sufficient for our goal of analyzing the overall stability and noise amplification of DDIM downscaling, without delving into spatial dependencies between adjacent pixels. In the following, we report for simplicity the results for the scalar case (i.e., $n=1$), and we leave the complete proofs for the general $n$-dimensional case to the appendix.
The theoretical derivation describing how the variance evolves through the diffusion process requires estimating the variance $\boldsymbol{v}_{t}$ at the $t$-th step based on $\boldsymbol{v}_{t-\Delta t}$ from the previous step. The following Proposition establishes a linear relationship between the two quantities.
@@
$\boldsymbol{F}_{t}:=\left(\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}+c_{t-\Delta t}\boldsymbol{J}_{t-\Delta t}\right)^{2},$
$\boldsymbol{g}_{t}:=c_{t-\Delta t}^{2}\boldsymbol{v}(\boldsymbol{r}_{t}),$
and $\boldsymbol{r}_{t}$ is a random variable independent on $\boldsymbol{x}_{t}$.
@@
Linearizing $\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})$ around $\boldsymbol{m}_{t-\Delta t}:=\mathbb{E}\left[\boldsymbol{x}_{t-\Delta t}\right]$ via the first order Taylor decomposition yelds:
$\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})\approx\boldsymbol{E}_{t-\Delta t}+\boldsymbol{J}_{t-\Delta t}\left(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t}\right)+\boldsymbol{r}_{t},$ (12)
where we defined $\boldsymbol{E}_{t-\Delta t}:=\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})$, $\boldsymbol{J}_{t-\Delta t}:=\nabla_{\boldsymbol{m}_{t-\Delta t}}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{m}_{t-\Delta t})$, and $\boldsymbol{r}_{t}$ is a remainder term, independent on $\boldsymbol{x}_{t-\Delta t}$. Plugging this expression into (4) leads to
$\boldsymbol{x}_{t}$ $=\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}\boldsymbol{x}_{t-\Delta t}+c_{t-\Delta t}\boldsymbol{\epsilon}_{\Theta}^{t-\Delta t}(\boldsymbol{x}_{t-\Delta t})$
$\approx\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}\boldsymbol{x}_{t-\Delta t}+c_{t-\Delta t}\left(\boldsymbol{E}_{t-\Delta t}+\boldsymbol{J}_{t-\Delta t}\left(\boldsymbol{x}_{t-\Delta t}-\boldsymbol{m}_{t-\Delta t}\right)+\boldsymbol{r}_{t}\right)$
$=\left(\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}+c_{t-\Delta t}\boldsymbol{J}_{t-\Delta t}\right)\boldsymbol{x}_{t-\Delta t}+c_{t-\Delta t}\boldsymbol{r}_{t}+c_{t-\Delta t}\boldsymbol{E}_{t-\Delta t}-c_{t-\Delta t}\boldsymbol{m}_{t-\Delta t}\boldsymbol{J}_{t-\Delta t}.$
Computing the variance on both side of the above equation, using the bilinearity of the variance, and ignoring terms that are constant in $\boldsymbol{x}_{t-1}$ as their variance is zero, yields:
$\boldsymbol{v}(\boldsymbol{x}_{t})\approx\left(\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}+c_{t-\Delta t}\boldsymbol{J}_{t-\Delta t}\right)^{2}\boldsymbol{v}(\boldsymbol{x}_{t-\Delta t})+c_{t-\Delta t}^{2}\boldsymbol{v}(\boldsymbol{r}_{t}),$
which corresponds to (11) by setting $\boldsymbol{F}_{t}:=\left(\sqrt{\frac{\alpha_{t}}{\alpha_{t-\Delta t}}}+c_{t-\Delta t}\boldsymbol{J}_{t-\Delta t}\right)^{2}$ and $\boldsymbol{g}_{t}:=c_{t-\Delta t}^{2}\boldsymbol{v}(\boldsymbol{r}_{t})$. ∎
Equation (11) represents a first-order linear recurrence relation of dimension $n$ with non-constant coefficients, for which a closed-form expression can be derived to compute $\boldsymbol{v}_{T}$, the variance of the generated data, as summarized in the following Theorem.
@@
Let $\boldsymbol{F}_{t}$ and $\boldsymbol{g}_{t}$ defined as in Proposition 3.1, and let $\Delta t$ be a fixed step-size for DDIM, so that any timestep $t$ can be written as $i\Delta t$ for $i=0,\dots,N$. Note that $\boldsymbol{v}_{0}=1$, as $\boldsymbol{x}_{0}\sim\mathcal{N}(0,1)$. Then, defining $\boldsymbol{F}_{(N+1)\Delta t}=1$ to simplify the notation, it holds:
$\boldsymbol{v}_{T}\approx\left(\prod_{i=1}^{N}\boldsymbol{F}_{i\Delta t}\right)+\sum_{i=1}^{N}\left[\left(\prod_{k=i+1}^{N+1}\boldsymbol{F}_{k\Delta t}\right)\boldsymbol{g}_{i\Delta t}\right].$ (13)
@@
The proof of this Theorem is mainly a recursive application of Proposition 3.1 to compute $\boldsymbol{v}_{T}$ from $\boldsymbol{v}_{0}$. Indeed, we recall that:
$\boldsymbol{v}_{t}\approx\boldsymbol{F}_{t}\boldsymbol{v}_{t-\Delta t}+\boldsymbol{g}_{t}.$
Since, as already observed, $\boldsymbol{v}_{0}=1$, then:
$\boldsymbol{v}_{\Delta t}\approx\boldsymbol{F}_{\Delta t}+\boldsymbol{g}_{\Delta t},$
$\boldsymbol{v}_{2\Delta t}\approx\boldsymbol{F}_{2\Delta t}\boldsymbol{F}_{\Delta t}+\boldsymbol{F}_{2\Delta t}\boldsymbol{g}_{\Delta t}+\boldsymbol{g}_{2\Delta t},$
$\boldsymbol{v}_{3\Delta t}\approx\boldsymbol{F}_{3\Delta t}\boldsymbol{F}_{2\Delta t}\boldsymbol{F}_{\Delta t}+\boldsymbol{F}_{3\Delta t}\boldsymbol{F}_{2\Delta t}\boldsymbol{g}_{\Delta t}+\boldsymbol{F}_{3\Delta t}\boldsymbol{g}_{2\Delta t}+\boldsymbol{g}_{3\Delta t},$
$\dots$
$\boldsymbol{v}_{N\Delta t}\approx\left(\prod_{i=1}^{N}\boldsymbol{F}_{i\Delta t}\right)+\sum_{i=1}^{N}\left(\prod_{k=i+1}^{N+1}\boldsymbol{F}_{k\Delta t}\right)\boldsymbol{g}_{i\Delta t},$
concluding the proof. ∎
Notably, Theorem 3.2 not only enables the explicit computation of the variance $\boldsymbol{v}_{T}$, but also highlights its dependence on the number of diffusion timesteps. A key result in this regard is presented in the following Theorem.
Theorem 3.3.
Let $\boldsymbol{F}_{t}$ and $\boldsymbol{g}_{t}$ defined as in Proposition 3.1, and let $\Delta t$ be the discrete step-size for DDIM, so that any timestep $t$ can be written as $i\Delta t$ for $i=0,\dots,N$, and $N=\frac{T}{\Delta t}$ Then $\{\boldsymbol{v}_{N\Delta t}\}_{N\in\mathbb{N}}$ is a monotonically increasing sequence and:
$\lim_{N\to\infty}\boldsymbol{v}_{N\Delta t}=e^{-\int_{0}^{T}\lambda(s)ds}+\int_{0}^{T}e^{-\int_{0}^{s}\lambda(\kappa)d\kappa}\gamma(s)ds,$ (14)
where:
$\begin{split}&\lambda(t)=-\partial_{t}\log\alpha(t-\Delta t)+\boldsymbol{J}_{t-\Delta t}\alpha^{\prime}(t-\Delta t)\left(\frac{\sqrt{1-\alpha(t-\Delta t)}}{\alpha(t-\Delta t)}+\frac{1}{\sqrt{1-\alpha(t-\Delta t)}}\right),\\ &\gamma(t)=\left(\frac{\alpha^{\prime}(t-\Delta t)\sqrt{1-\alpha(t-\Delta t)}}{2\alpha(t-\Delta t)}+\frac{\alpha^{\prime}(t-\Delta t)}{2\sqrt{1-\alpha(t-\Delta t)}}\right)\boldsymbol{v}(\boldsymbol{r}_{t})\end{split}$
Sketch of the proof.
We report here only the main ideas of the derivation, while the complete proof is provided in Appendix A.1. A first-order Taylor expansion of $\sqrt{\alpha(t)}$ and $\sqrt{1-\alpha(t)}$ around $t-\Delta t$ shows that:
$\boldsymbol{F}_{t}=1+\lambda_{t-\Delta t}\Delta t+O(\Delta t^{2})$
$\boldsymbol{g}_{t}=\gamma_{t-\Delta t}\Delta t+O(\Delta t^{2}),$
where $\lambda_{t-\Delta t}$ and $\gamma_{t-\Delta t}$ depend on $\alpha^{\prime}(t-\Delta t)$ and the local Jacobian $\boldsymbol{J}_{t-\Delta t}$.
Substituting in the expression of Theorem 3.2 leads to a product–sum expression whose discrete factors $\prod_{k=1}^{T}\boldsymbol{F}_{k}$ converge monotonically from below to $\exp\!\big(\int_{0}^{T}\lambda(s)\,ds\big)$ as $\Delta t\to 0$. The variance sequence $\{\boldsymbol{v}_{N\Delta t}\}_{N\in\mathbb{N}}$ therefore satisfies
$\boldsymbol{v}_{N\Delta t}\overset{N\to\infty}{\longrightarrow}e^{-\int_{0}^{T}\lambda(s)\,ds}+\int_{0}^{T}e^{-\int_{0}^{s}\lambda(\kappa)\,d\kappa}\,\gamma(s)\,ds,$
completing the proof. ∎
Theorem 3.3 reveals a compelling property regarding the relationship between the variance of the diffusion process at the final step, $\boldsymbol{v}_{T}$, and the number of diffusion steps performed. Specifically, it shows that the variance increases monotonically with the number of steps $N$, demonstrating that it is possible to calibrate the variance of the generated data by adjusting the number of diffusion steps, as further discussed in the next section. Moreover, (14) indicates that $\boldsymbol{v}_{T}$ cannot grow indefinitely, but instead converges to a finite limit determined by the schedule $\{\alpha_{t}\}_{t\in[0,T]}$ and by the neural network $\epsilon_{\Theta}^{t}$ through the Jacobian $\boldsymbol{J}_{t}$. To the best of our knowledge, this is the first work that theoretically proves that the variance expressed by diffusion models can be calibrated by tuning the number of diffusion steps. In the following section, we empirically demonstrate that this property holds in practice by testing variance calibration on conditioned diffusion models applied to ERA5-to-CERRA downscaling.
@@ 4.2 Training on the full domain
The general structure of our diffusion model follows the framework established in [Merizzi2024]. The denoiser is a residual U-Net [ronneberger2015, 7780459] implemented in TensorFlow/Keras, designed as a symmetric encoder–decoder network with skip connections and residual convolutional blocks using Swish activations and layer normalization. The architecture performs three stages of downsampling through average pooling and three corresponding upsampling stages via bilinear interpolation. Each stage (both in the encoder, decoder, and latent bottleneck) contains three residual blocks, each composed of two 3×3 convolutions and a skip connection to preserve spatial detail. The model contains approximately 22 million learnable parameters, predicting the noise component at each diffusion step.
Training is conducted by uniformly sampling signal-to-noise ratios from the sinusoidal noise scheduling functions, optimizing the mean absolute error (MAE) between the sampled noise and the predicted noise.
@@
Recent findings [Chen2023, Hoogeboom2023] have demonstrated that cosine noise schedules can reduce the effective noise added by the model as image size increases. This occurs because redundancy in data, such as correlations among nearby pixels, typically increases with resolution, and the independently added noise becomes easier to denoise in larger images. Consequently, noise schedules optimized for smaller resolutions may not perform as effectively at higher resolutions. To address this issue, we introduce a constant scaling factor $\lambda>1$ to reduce the signal rate in the noise schedule. For an input $\boldsymbol{x}_{T}\sim p_{gt}(\boldsymbol{x}_{T})$ and a scaling factor $\lambda$, the noise schedule eq. 1 is modified as follows:
@@
Experimentally, we tested several values of $\lambda\in[1,4]$ and observed that $\lambda=3$ provided the best trade-off between reconstruction fidelity and training stability, achieving the lowest MSE with respect to the deterministic CERRA ground truth. Full ablation at multiple $\lambda$ values and full resolution would be computationally prohibitive; therefore, smaller-scale experiments were used to guide the choice, keeping the spatial resolution constant as it is the key factor influencing this effect. The selected value should thus be regarded as empirically determined. This adjustment proved crucial for large-scale ensemble diffusion applications, improving convergence and overall stability. Notably, applying the same scaling to the low-resolution conditioning signal further stabilized the training process.
While the core U-Net architecture remained largely unchanged, scaling to larger domains required increasing the number of residual blocks in the bottleneck. This adjustment improved performance when generating larger images [Hoogeboom2023]. The model’s performance across the increased domain remained comparable to that observed at smaller resolutions, with the bottleneck modification enhancing its ability to process higher-resolution inputs effectively.
The diffusion model is trained on the ERA5 to CERRA task for wind speed on the full domain, utilizing data from 1985 to 2010 for training. The training was performed on a single compute node equipped with three NVIDIA A100 GPUs (64 GB VRAM each), using a batch separation (mirroring) strategy to distribute the workload across devices. The model was trained for 400 epochs with 500 steps per epoch, corresponding to a total training time of approximately 42 hours. Each epoch required about 300 seconds to complete, and the batch size was set to three, with one sample processed per GPU. AdamW was used as the optimizer, with an initial training phase of 200 epochs using a learning rate of $1\times 10^{-4}$ and a weight decay of $1\times 10^{-5}$, followed by a fine-tuning phase of 100 epochs with a reduced learning rate of $1\times 10^{-5}$ and a weight decay of $1\times 10^{-6}$. The model was trained to minimize the mean absolute error (MAE) between the predicted and true noise. During inference, the memory constraints are relaxed, and the model can be executed on GPUs with as little as 16 GB of VRAM.
4.3 Implementation Details
The denoiser is a residual U-Net implemented in TensorFlow/Keras, featuring symmetric encoder–decoder paths with skip connections and residual convolutional blocks using Swish activations and layer normalization. Downsampling is done via average pooling, upsampling via bilinear interpolation, and the sinusoidal embedding of the diffusion timestep is concatenated to the input features
Please add model size, training duration, GPU hours, and inference latency vs. steps $N$ on the full grid (1069×1069), plus memory use. This will help weigh calibration quality against runtime; your hardware/optimizer details are noted but not quantified.
@@ 5 Experiments
Our primary experimental objective is to evaluate the behavior of ensemble variance while changing the timestep $\Delta t$. While a comprehensive analysis of the ensemble diffusion model’s performance across the entire CERRA domain is beyond the scope of this paper, we ensure that the model operates within a realm of satisfactory performance. To achieve this, we begin—similar to [Merizzi2024]—by comparing our model’s performance against bilinearly interpolated ERA5 data and a residual U-Net model trained for the same task. The results, obtained by computing the Mean Squared Error (MSE) and SSIM between the high-resolution output produced by the model and the original CERRA, are shown in Table 1.
@@
MSE $(\downarrow)$ $3.73\text{e-03}$ $2.50\text{e-04}$ $2.54\text{e-04}$
SSIM $(\uparrow)$ $0.751$ $0.892$ $0.923$